OpenStudy (anonymous):

Use the elimination method to solve the system for x and y. 6x+y=51 x-2y=2 Select one: a. (8, 3) b. (3, -8) c. (3, 8) d. (-8, 3)

OpenStudy (anonymous):

6x+y=51 (multiply by 2) =>12x+2y=102 + x-2y=2 ________________________ 13x =104 x=104/13 x=8 substitute x=8 into 1: 6(8)+y=51 48+y=51 y=51-48 y=3 Therefore it is a. (8,3)

OpenStudy (anonymous):

ok thanks!(: i appreciate it